Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
boldrini, mantovani, ruggi, nardecchia - 15:55 Friday 24 September 2021 (53264) Print this report
ISC shift: green laser glitch investigation, ITF adjustment, TCS tuning

The goal of the shift was to complete the recovery of the ITF and continue with the tuning of TCS.

At the beginning of the shift we waited for ALS team to complete their intervention on the green laser (53262). The effectiveness of this intervention is yet to be determined by analyzing the data, even if the situation seems to be improved.  We will leave the green locked for the whole weekend.

Two other attempts were interrupted by a failure with the CARM hand-off to the IR signal. We tried the solution proposed in (53253) and opened the drift control of the terminal benches and the problem did not present again, but being it so sporadic we cannot tell whether this solution is making a real difference. This procedure will have to be integrated in automation.

Since we encountered the already observed problems at the various steps of the lock acquisition performed with the goto functions, namely the 55 Hz possibly given by low DARM gain or high PRCL gain (may be the DARM dem phase tuning could be tuned), we decided to change the 480mW step to 600mW step (which is more similar in Hz to what we had before the increase of intracavity power thanks to the RH). The last function before carm null is now called goto_600mW_HP(). No other changes have been made to the functions concerning gains or phases.

The Carm null has been reached without the wait of 5mins at this step (this was done in order to do not have the CARM offset drift fast and in un-controlled way due to the high thermal effects since the CARM error signal is generated by B7 and B8 DC) .

From the behavior of the DIFFp signals a clear mistuning of the demodulation phase was present (expecially on the horizontal direction). The first task that will be indeed done in the next ISC shift will be the DiffP phase tuning.

Comments to this report:
mantovani - 16:32 Friday 24 September 2021 (53267) Print this report

As it is visible from the B4 6MHz I signal the 600mW is more or less equivalent (slightly lower) to the old 480mW step.

Images attached to this comment
mantovani - 17:05 Friday 24 September 2021 (53269) Print this report

the mistuining of the Diffp demodulation phases is evident both in ty and tx @ Carm null. see attached plots

Images attached to this comment
Search Help